OpenStudy (221emily):

I need Help with Algebra 2 I will fail if i don't get this done. I will fan and medal anyone who gives me some serious help.

OpenStudy (sagewilson):

Give me one min, you don't have to fan or medal me, just let me finish up something iim working on.

OpenStudy (seanstewart16):

lets see what my iq gets me

OpenStudy (seanstewart16):

what are we doing here ?

OpenStudy (221emily):

A factory can produce two products, x and y, with a profit approximated by \[p=14x+22y-900\] The production of y must exceed the production of x by at least 100 units. Moreover, production levels are limited by the formula \[x+2y \le1400\] a. identify the vertices of the feasible region b. what production levels yield the maximum profit, and what is the maximum profit?

OpenStudy (221emily):

I would like to be walked through it. If someone could please specify, like this is step one, you do this and you get this and then step two and so on

OpenStudy (hamde):

y exceed x by 100 then we can say that y-x=100 ....... a the inequality gives us two propabilites first x+2y=1400 .......b or second x+2y>1400......c from a,b we get that y=500 by substituting x=600 so we can get the profit p=14*600+22*500-900 but from a,c we get y>500 , x>600 then the max profit will be p>14*600+22*500-900 is that clear emi i hope so .